OpenStudy (anonymous):

any help again? just asking for double check. :) Two kinds of coffee are blended to produce a mixture of 25 kilograms worth $52 a kilogram. The costs of the two kinds of coffee are $40 and $56 a kilogram. Find how many kilograms of each kind are blended. last probleeem! NEED HELP! :)

OpenStudy (anonymous):

need solution. :')

OpenStudy (anonymous):

how did you attempt to set up your problem?

OpenStudy (anonymous):

x + y = 25 40x + 56y = 52

OpenStudy (anonymous):

solve that system

OpenStudy (anonymous):

how can i solve it if two variables are in the same equation?

OpenStudy (anonymous):

what math class are you in? have you done systems of equations yet?

OpenStudy (amistre64):

40x + 56y = 52(25) right?

OpenStudy (anonymous):

oops yeah

OpenStudy (anonymous):

is the answer 6.25@40?

OpenStudy (amistre64):

lol ... when I suggested you dbl chk my results; I meant plug them into the equations and see if the results are valid ;)

OpenStudy (anonymous):

i did that already. i just asked to make sure cuz this one's pretty hard btw, thanks tho. :)

OpenStudy (amistre64):

:) when I plug this into my calculator: 40(6.25) + 56(18.75) = 52(25) it tells me its good

OpenStudy (anonymous):

all right. :)